Jack spends his whole budget and purchases 8 units of hamburger and 5 units of cake. The price of hamburger is twice the price of cake. The consumer’s income doubles and the price of cake doubles, but the price of hamburger stays the same. If the consumer continues to buy 5 units of cake, what is the largest number of hamburger that he can afford?
1
Expert's answer
2016-01-19T08:28:41-0500
Jack purchases 8 units of hamburger Qh = 8 and 5 units of cake Qc = 5, Pc = 2Ph, I = 8Ph + 5Pc Inew = 2I, Pc(new) = 2Pc, so Pc(new) = 4Ph
If the consumer continues to buy 5 units of cake, then the largest number of hamburger that he can afford will be:
I = 8Ph + 5Pc = 8Ph + 10Ph = 18Ph
Inew = 2I = 5Pc(new) + Qh*Ph = 20Ph + Qh*Ph
18Ph*2 = 20Ph + Qh*Ph
16Ph = Qh*Ph
Qh = 16 - is the largest number of hamburger that Jack can afford after increase in income.